OpenStudy (anonymous):

-4n=28n-2 A) 1/16 b)-2/32 c) no solution d) all real numbers

OpenStudy (anonymous):

-4n = 28n - 2 -32n = -2 n = 2/32 Simplify the fraction by taking out 2. n = 2/32 n = 1/16